Dear math teacher,
I am having difficulties solving for n in the following problem:
nC12 = nC8
n!/(n-12)!12! = n!/(n-8)!8!
n(n-1)...(n-11)(n-12)!/(n-12)!12! = n(n-1)...(n-7)(n-8)!/(n-8)!8!
(n-8)(n-9)(n-10)(n-11)/12! = 1/8!
n^4-21n^3+110n^2-17n^3+357n^2-1870n+72n^2-1512n+7920 / 12! = 1/8!
Would you please let me know whether I am on the right track? My solving steps seem convoluted like a kidney tubule! Please let me know where I made an error.
Thank you very much.
Yours,
I.

Answer by Theo(13342)

you're on the right track, but you needed to do some other type thinking before getting to the solution, which is a rule that i will explain to you after.
the equation is:
nC12 = nC8
this formula translates to:
n!/(12!*(n-12)!) = n!/(8!*(n-8)!)
cross multiply to get:
n! * (8!*(n-8)!) = n! * (12!*(n-12)!)
divide both sides of this equation by n! to get:
8! * (n-8)! = 12! * (n-12)!
divide both sides of this equation by 8! and divide both sides of this equation by (n-8)! to get:
(n-8)! / (n-12)! = 12!/8!
12! is equal to 12 * 11 * 10 * 9
(n-8)! / (-12)! is equal to (n-8)*(n-9)*(n-10)*(n-11)*(n-12)! / (n-12)! which is equal to:
(n-8)*(n-9)*(n-10)*(n-11)
our equation becomes:
(n-8)*(n-9)*(n-10)*(n-11) = 12*11*10*9
the left side of this equation will be equal to the right side of this equation if:
n-8 = 12
n-9 = 11
n-10 = 10
n-11 = 9
all of these equation would be true if n = 20.
that's your answer.
n = 20.
20C12 = 125970
20C8 = 125970
they're the same.
note that 12 + 8 = 20
that's the rule.
you can use it with any problems of this type to immediately get the answer.
an example:
nC13 = nC22
13 + 22 = 35
your answer is n = 35
35C13 = 1476337800
35C22 = 1476337800
The basic properties of combinations allows this to happen.
that property is that nCx = nC(n-x)
from that formula, you get x + (n-x) = n
a simpler example:
take n = 7
your values for each of the possible combinations are:
7C0 = 1
7C1 = 7
7C2 = 21
7C3 = 35
7C4 = 35
7C5 = 21
7C6 = 7
7C7 = 1
you can see that 7Cx = 7C(7-x)
this makes 7C0 = 7C7
this makes 7C1 = 7C6
this makes 7C2 = 7C5
etc.
Add the x and the (n-x) together and you get 7 every time.
your problem as nC12 = nC8
add 12 and 8 together and you get 20.